# Learning Fine Pinch-Grasp Skills using Tactile Sensing from Real Demonstration Data

### v1

@@ Abstract
This work develops a data-efficient learning from demonstration framework which exploits the use of rich tactile sensing and achieves fine dexterous bimanual manipulation. Specifically, we formulated a convolutional autoencoder network that can effectively extract and encode high-dimensional tactile information. Further, we developed a behaviour cloning network that can learn human-like sensorimotor skills demonstrated directly on the robot hardware in the task space by fusing both proprioceptive and tactile feedback. Our comparison study with the baseline method revealed the effectiveness of the contact information, which enabled successful extraction and replication of the demonstrated motor skills. Extensive experiments on real dual-arm robots demonstrated the robustness and effectiveness of the fine pinch grasp policy directly learned from one-shot demonstration, including grasping of the same object with different initial poses, generalizing to ten unseen new objects, robust and firm grasping against external pushes, as well as contact-aware and reactive re-grasping in case of dropping objects under very large perturbations. Moreover, the saliency map method is employed to describe the weight distribution across various modalities during pinch grasping. The video is available online at: https://youtu.be/4Pg29bUBKqs.
@@ I INTRODUCTION
Tactile sensing plays a vital role in capturing detailed information about contact surfaces, including the distribution of contact forces and their variations during force-sensitive tasks – which is indispensable for achieving dexterous handling of lightweight objects with irregular surfaces, shapes, and deformable properties. Especially during close-range interaction between hands and objects, visual occlusion restricts the ability to perceive detailed information of the contact surfaces, during which tactile sensors become valuable for providing essential information of these unseeable surfaces. Integrating tactile sensing into motor learning of dexterous grasping can enhance the rich and precise sensing of surface contacts and interaction dynamics, provide irreplaceable and direct feedback when manipulating objects [4], and enable more robust and precise manipulation tasks. It is crucial to explore how robots can leverage this information to achieve human-comparable dexterous manipulation abilities.
@@
The canonical hardware for robot manipulation incorporates Force/Torque sensors that can only measure the 6-degree-of-freedom (DoF) wrench at each end-effector. Soft optical-based tactile sensors can provide abundant and discriminative contact information by quantifying the deformation of the soft materials using a camera system [5]. Currently, several soft tactile sensors have been developed, including TacTip [6], DigiTac [7], Gelsight [8], and DIGIT [9]. However, how to use high-dimensional data from tactile sensors for robot dexterous grasping remains open research.
@@
The complex and non-trivial deformation of soft tactile sensors during dexterous grasping tasks presents a considerable challenge. Human can deal with soft contacts, quuckly adapt to new tasks, and produce skills of dual-arm coordination for manipulating objects. Learning from Demonstration (LfD) offers an intuitive, efficient method for acquiring human skills through synchronized tactile information, encoding rich state-action mapping and enabling robots to learn human sensorimotor skills while responding to tactile and proprioceptive feedback. In addition, the common issue of accumulating compounding errors during dexterous manipulation task execution in LfD can be mitigated by utilizing rich tactile information as feedback. The challenge involves effectively extracting features from sensory data and integrating them with proprioceptive states for sample-efficient human dexterous manipulation behavior learning.
This work is motivated to develop an effective LfD framework that leverages rich tactile sensing to learn dexterous sensorimotor skills. Our approach focuses on achieving one-shot LfD of fine pinch grasp, using high-dimensional contact information from tactile sensors and a limited amount of real data. The contributions are summarized as follows:
  • A novel feature extraction approach to encapsulate essential features from tactile sensing data, which are then fused with robot proprioceptive states and tactile image difference, thus resulting in a low-dimensional latent space representation that significantly enhances the learning process of fine grasping skills.
  • An effective LfD framework that integrates tactile sensory input and robot proprioceptive state, which enables the robot to efficiently acquire feedback-driven dexterous grasping skills through a single demonstration.
@@
The proposed framework is validated by pinch grasp tasks on a dual-arm setup equipped with TacTips sensors [6] and has achieved the successful retrieval of a small, cylindrical object on a table using one-shot demonstration. Our experimental results show that the policy, learned from one-shot human demonstration data, can achieve stable grasping of unseen objects with different diameters, masses, and materials.
@@
Furthermore, the robustness of the framework against external disturbances has been validated, with the learned policy demonstrating stable grasping under external disturbance, as well as the capacity to autonomously execute successful re-grasping in case of a large external force that pushes off the object. We applied saliency map analysis [10] and revealed how the learned policy uses different sensory modalities in a variable way throughout the dexterous pinch grasp process, and demonstrate the capability and effectiveness of our proposed network to efficiently learn features of the high-dimensional data and autonomously segment the long-horizon data into several distinct fine-skills for execution.
@@ II Related Works
During robotic grasping, tactile sensors can provide rich contact information which is not easily accessed via visual information, thereby playing a crucial role in enhancing the dexterous grasping capabilities [11]. Prior research on robotic pinch grasp has primarily focused on either force analysis and planning to achieve force closure [12] or the development of specialized grippers [13]. Soft deformable tactile sensors have the ability to perform contact-rich interactions with the environment and manipulate delicate objects safely [14]. With optical-based tactile sensors, the orientation of the contact surface can be inferred from the tactile image, enabling stabilization of the pinch grasp by rolling the sensor on the contact surface and applying desired grasping forces [15]. The study in [16] proposed a novel tactile sensor capable of measuring and localizing distributed forces that enables the robot hand to grasp deformable soft objects.
One open question is how to extract useful information from high-dimensional tactile images. The works in [17] estimate 6D contact wrenches from tactile images and the estimated wrenches that can be used as feedback to the grasping controllers within the classical control theory. Deep neural networks can also be used to process tactile images. The works in [18] show that contact poses can also be detected from tactile images, which was then combined with goal-driven methods to achieve non-prehensile object stable pushing. The works in [19] introduce Autoencoder networks [20] to compress the high-dimensional tactile images into low-dimensional latent vectors which can be used for several down-stream tasks, such as object classification.
@@ III Methods
(a) The Convolutional Autoencoder network for extracting tactile information into reduced dimensional latent vector.
(b) The Behaviour Cloning network that fuses the tactile sensing, normal contact forces and proprioceptive information.
Figure 3: The structure of Convolution Autoencoder netowrk and Behavior Cloning network.
@@ III-A System Overview
Teleoperation through a physical robot is a viable approach for generating real demonstration data that can be executed on a physical system, and it was shown to be effective to perform fine dexterous grasping [24]. As shown in Fig. 2, the overall architecture incorporates a teleoperation system for the collection of human demonstration data, and a dual-arm setup for executing pinch grasp tasks. The teleoperation system consists of two haptic devices (Force Dimension Sigma 7) for human operators to control the dual-arm robot [25]. The dual-arm robot system includes two Franka Emika Panda arms each with a TacTip [6] installed on the end-effector of each arm. The Tactips capture contact information between end-effector and objects as 2D tactile images. Task-Space Sequential Equilibrium and Inverse Kinematics Optimization (SEIKO) runs in the backend to guarantee the guarantee of physical constraints and safety of the dual-arm robot [26].
The Learning from Demonstration (LfD) framework (see Fig. 3) is composed of two distinct networks: 1) a Convolutional AutoEncoder (CAE) network to extract the latent features from tactile images; 2) a Behaviour Cloning (BC) network to learn the policy of dexterous dual-arm grasping with tactile sensing from human demonstrations.
@@ III-C Tactile Feature Extraction
The Tactip used in this work is an optical tactile sensor with a soft hemispherical tip, which was 3D-printed in one piece combining an elastic skin with 330 rigid white markers (pins) [6]. When the soft tip deforms during contact with objects, the white pins start to move away from their initial positions. The displacement of these pins reflects the complex deformation of the soft surface. An inner camera captures and projects the displacement to an array of white pins on a black background in the image plane, as shown in Fig. 3a. Raw tactile RGB images are firstly resized to 256$\times$256 pixels using linear interpolation and converted to grayscale images, which are then cropped using a circle mask and converted to binary images by thresholding. A median filter is applied to denoise the binary images.
We propose to use a self-supervised learning method – convolutional autoencoder network to extract robust features that can represent the contact properties from the preprocessed tactile images. Eight convolutional layers are used in the CAE network to extract the spatial information represented by the displacement of the pins. The structure of CAE network is shown in Fig. 3a. The CAE network consists of an encoder and a decoder, formulated as follows:
@@ III-D Behavior Cloning Network
We propose and design a novel BC network to learn the behaviors of coordinated manipulation skills of bimanual grasping from human demonstration data. Dexterous bimanual grasping skills can be considered into two categories: (1) adaptive interaction with different objects, and (2) dual-arm motion coordination. To capture these skills, we have designed the input to our network to include encoded tactile feature maps, tactile image differences, and the robot’s proprioceptive state. The encoded feature maps and tactile image differences capture the human-object interaction skills. The robot’s proprioceptive state, on the other hand, offers insights into the coordination of movements between both arms. These inputs collectively serve to reflect the complexity and adaptability of dexterous grasping skills.
@@
Using such data of multiple modalities as input to train a network requires a well-crafted embedding structure. A common way of fusing a 2D feature map and a 1D feature vector is to flatten the 2D feature map into a 1D vector and concatenate the flattened vector and the 1D feature vector. However, we found that the flattening projection results in the loss of spatial correlation of tactile information. In this work, we specifically tile the proprioceptive state of robots and the tactile image difference to match the dimension of the tactile feature maps, so as to keep the spatial information of the encoded tactile feature maps.
@@
We then concatenate the tactile feature maps, the tiled proprioceptive state maps, and the tactile image difference on each feature channel, as shown in Fig. 3b. The convolutional layers in the BC network first filter the input feature maps (46$\times$16$\times$16) to a feature map (1$\times$8$\times$8), which is then flattened and fed into a fully connected network (FCN). The FCN network outputs a vector $\hat{\bm{a}}\in\mathbb{R}^{12}$ as the predicted Cartesian pose commands of the two arms, including 3D position and 3D orientation for each arm.
@@ IV-A Experimental Setup and Data Collection
We validate the performance of LfD with tactile sensing for robot dexterous manipulation on the challenging tasks: object retrieval from the desk using dual-arm pinch grasp. We have designed a comparative study involving two different configurations to show that our method can outperform the vanilla BC baseline qualitatively and quantitatively. During dexterous grasping, external vision can be easily occluded by the end-effector, potentially leading to inaccurate object estimation. Therefore, our experiments operate without using external visual sensors. By default, the starting position of the object lies between two robot hands, and the whole demonstration is represented in the task space. This assumption allows us to have an interface to connect with a 6D pose estimation from an external camera (e.g., as in [27]) and then position the initial poses of two hands around the target object before starting our control policy.
@@
We collected two demonstrations for each task. The human demonstration dataset collected in the grasping task includes three main components: the Cartesian commands, the proprioceptive states, and the tactile feedback (i.e., tactile images provided by the TacTip sensors). The Cartesian commands and the proprioceptive states of the two arms are collected at a frequency of 1000 Hz. Two Tactips record the tactile image pairs at a frequency of 60 Hz. For each demonstration, about 1000 tactile images are recorded. Before using the collected dataset to train the networks, several pre-processing methods are used to process the raw data. The proprioceptive states of the two arms and the tactile images, collected at different sampling rates, are synchronized using a linear interpolation method to align their timestamps. A median filter is then applied to smooth the Cartesian commands $a_{t}$, i.e., the 6D poses of two end-effectors. For raw tactile images, the structural similarity index measure (SSIM) [28] is used to quantify the difference between the current frame and the original frame.
IV-B Design of Validation Tasks
IV-B1 Learning grasping vial
@@
The human demonstrator performs teleoperation of dual-arm robots to grasp a plastic vial (a test tube with $\Phi=15.65$mm) that is horizontally placed on the table. A Behaviour Cloning (BC) network is trained using the garthered demonstration data, and the trained policy is tested on dual-arm robots to validate its generalisation on unseen initial poses. During the evaluative phase, we positioned the test tube between the end-effectors to evaluate the performance of the learned policy given variations in the starting position, specifically alterations of up to ±$\pm 20$ degrees and displacements of up to ±$\pm 2$ centimetres in the objects’ locations.
IV-B2 Generalization to unseen objects
@@
IV-B3 Robustness against external disturbance
@@ IV-B4 Re-grasping capability
IV-C Network Evaluation
Our proposed model is developed using PyTorch [29]. For the training of the Convolutional AutoEncoder (CAE), a dataset comprising 1500 tactile images obtained from the TacTip sensor during the demonstration was utilized as a demonstration. A representation of reconstruction performance is shown on the validation set is shown in Fig. 3a. The trained CAE exhibits a satisfactory reconstruction quality, with a Mean Squared Error (MSE) loss of 0.015 and a Structural Similarity Index Measure (SSIM) of 0.934. The model training process, which involved 100 iterations, was completed in approximately two hours using an NVIDIA 1080 GPU. In the case of the Behaviour Cloning (BC) network, the model was trained for 1000 iterations and the training process takes approximately 5 minutes.
@@ IV-C Network Evaluation
(b) Unsuccessful grasping of the baseline FCN policy trained with only the prioperceptive information of end-effectors’ pose and positions.
Figure 5: Comparison study using the baseline method (failed) in real-world robots.

### v2

# Learning Fine Pinch-Grasp Skills using Tactile Sensing from A Few Real-world Demonstrations

@@ Abstract
Imitation learning for robot dexterous manipulation, especially with a real robot setup, typically requires a large number of demonstrations. In this paper, we present a data-efficient learning from demonstration framework which exploits the use of rich tactile sensing data and achieves fine bimanual pinch grasping. Specifically, we employ a convolutional autoencoder network that can effectively extract and encode high-dimensional tactile information. Further, We develop a framework that achieves efficient multi-sensor fusion for imitation learning, allowing the robot to learn contact-aware sensorimotor skills from demonstrations. Our comparision study against the framework without using encoded tactile features highlighted the effectiveness of incorporating rich contact information, which enabled dexterous bimanual grasping with active contact searching. Extensive experiments demonstrated the robustness of the fine pinch grasp policy directly learned from few-shot demonstration, including grasping of the same object with different initial poses, generalizing to ten unseen new objects, robust and firm grasping against external pushes, as well as contact-aware and reactive re-grasping in case of dropping objects under very large perturbations. Furthermore, the saliency map analysis method is used to describe weight distribution across various modalities during pinch grasping, confirming the effectiveness of our framework at leveraging multimodal information. The video is available online at: https://youtu.be/BlzxGgiKfck.
@@ I INTRODUCTION
Tactile sensing plays a vital role in capturing detailed information about contact surfaces, including the distribution of contact forces and their variations during force-sensitive tasks – which is indispensable for achieving dexterous handling of lightweight objects with irregular surfaces, shapes, and deformable properties. Especially during close-range interaction between hands and objects, visual occlusion restricts the ability to perceive detailed information of the contact surfaces, during which tactile sensors become valuable for providing essential information of these unseeable surfaces. Integrating tactile sensing into motor learning of dexterous grasping can enhance the rich and precise sensing of surface contacts and interaction dynamics, provide irreplaceable and direct feedback when manipulating objects, and enable more robust and precise manipulation tasks [4, 5]. It is crucial to explore how robots can leverage this information to achieve dexterous manipulation abilities.
@@
The canonical hardware for robot manipulation incorporates Force/Torque sensors that can only measure the 6-degree-of-freedom (DoF) wrench at each end-effector. Soft optical-based tactile sensors can provide abundant and discriminative contact information by quantifying the deformation of the soft materials using a camera system [5]. Currently, several soft tactile sensors have been developed, including TacTip [6], DigiTac [7], Gelsight [8], and DIGIT [9]. However, how to use high-dimensional data from tactile sensors for robot contact-rich tasks remains open research.
@@
The complex and non-trivial deformation of soft tactile sensors during dexterous grasping tasks presents a considerable challenge. Humans can deal with soft contacts, quickly adapt to new tasks, and produce skills of dual-arm coordination for manipulating objects. Learning from Demonstration (LfD) offers an intuitive, efficient method for acquiring human skills through synchronized tactile information, encoding rich state-action mapping and enabling robots to learn human sensorimotor skills while responding to tactile and proprioceptive feedback. Additionally, the issue of errors accumulating in contact-rich tasks during Learning from Demonstration (LfD), due to the lack of direct feedback, can be addressed by incorporating rich tactile feedback in real-time. The challenge involves effectively fusing high-dimensional data with robot proprioceptive states for sample-efficient human dexterous manipulation behavior learning.
I-A Contribution
@@
In this work, we present a framework to teach bimanual, dexterous sensorimotor skills. To handle the complex dynamics involved with deformable tactile sensors, we employ behavior cloning (BC) to learn from human teleoperated demonstrations. A convolutional autoencoder (CAE) is trained in a self-supervised manner to extract essential features from the rich tactile data, which are then integrated with the robot’s proprioceptive state. This multimodal fusion enables the robot to efficiently acquire feedback-driven dexterous grasping skills through a few human demonstrations. The proposed framework is validated by pinch grasp tasks on a dual-arm setup equipped with TacTips sensors [6] and has achieved the successful retrieval of a small, cylindrical object on a table using few-shot demonstrations. Our experimental results show that the policy, learned from few-shot human demonstration data, can achieve stable grasping of unseen objects with different diameters, masses, and materials.
@@
Furthermore, the robustness of the framework against external disturbances has been validated, with the learned policy demonstrating stable grasping under external disturbance, as well as the capacity to autonomously execute successful re-grasping in case of a large external force that pushes off the object. We applied saliency map analysis [10] and revealed how the learned policy uses different sensory modalities in a variable way throughout the dexterous pinch grasp process. This analysis demonstrates the capability and effectiveness of our proposed network to efficiently use high-dimensional data and autonomously segment the long-horizon data into several distinct fine-skills for execution according to different contact situations.
@@ II Related Works
During robotic dexterous manipulation, tactile sensors can provide rich contact information which is not easily accessed via visual information, thereby playing a crucial role in enhancing the dexterous grasping capabilities [11]. Soft deformable tactile sensors can perform contact-rich interactions with the environment and manipulate delicate objects safely [12]. With optical-based tactile sensors, the orientation of the contact surface can be inferred from the tactile image, enabling stabilization of the pinch grasp by rolling the sensor on the contact surface and applying desired grasping forces [13]. The study in [14] explores utilizing tactile feedback to achieve bimanual tasks including bi-pushing, bi-reorienting, and bi-gathering, conducted within a simulated environment using reinforcement learning, and successfully achieves simulation-to-reality transfer. In contrast, our work focuses on using both deformable properties and rich surface contact information provided by the tactile sensor to achieve dexterous bimanual pinch grasping of small, fragile, and delicate objects.
One open question with high-dimensional tactile sensors is how to extract useful information from them. The works in [15] estimate 6D contact wrenches from tactile images and the estimated wrenches that can be used as feedback to the grasping controllers within the classical control theory. Deep neural networks can also be used to process tactile images. The works in [16] show that contact poses can also be detected from tactile images, which was then combined with goal-driven methods to achieve non-prehensile object stable pushing. The works in [17] introduce Autoencoder networks [18] to compress the high-dimensional tactile images into low-dimensional latent vectors which can be used for several down-stream tasks, such as object classification. In our work, we similarly employ an autoencoder to extract and encode tactile images. However, we extend its utility to achieve data-efficient learning from human demonstration by fusing the encoded tactile feature with robot proprioceptive information.
@@ III-A System Overview
Teleoperation through a physical robot is a viable approach for generating real demonstration data that can be executed on a physical system, and it was shown to be effective in performing fine dexterous grasping [22]. As shown in Fig. 2, the overall architecture incorporates a teleoperation system for the collection of human demonstration data and a dual-arm setup for executing pinch grasp tasks. The teleoperation system consists of two haptic devices (Force Dimension Sigma 7) for human operators to control the dual-arm robot [23]. The dual-arm robot system includes two Franka Emika Panda arms each with a TacTip [6] installed on the end-effector of each arm. The Tactips capture contact information between end-effector and objects as 2D tactile images. Task-Space Sequential Equilibrium and Inverse Kinematics Optimization (SEIKO) runs in the backend to guarantee the guarantee of physical constraints and safety of the dual-arm robot [24].
The Learning from Demonstration (LfD) framework (see Fig. 2) is composed of two distinct networks: 1) a Convolutional AutoEncoder (CAE) network to extract the latent features from tactile images; 2) a Behavior Cloning (BC) network to learn the policy of dexterous dual-arm grasping with tactile sensing from human demonstrations.
@@ III-C Tactile Feature Extraction
The Tactip used in this work is an optical tactile sensor with a soft hemispherical tip, which was 3D-printed in one piece combining an elastic skin with 330 rigid white markers (pins) [6]. When the soft tip deforms during contact with objects, the white pins start to move away from their initial positions. The displacement of these pins reflects the complex deformation of the soft surface. An inner camera captures and projects the displacement to an array of white pins on a black background in the image plane. Raw tactile RGB images are firstly resized to 256$\times$256 pixels using linear interpolation and converted to grayscale images, which are then cropped using a circle mask and converted to binary images by thresholding. A median filter is applied to denoise the binary images.
We propose to use a self-supervised learning method – convolutional autoencoder network to extract robust features that can represent the contact properties from the preprocessed tactile images. Eight convolutional layers are used in the CAE network to extract the spatial information represented by the displacement of the pins. The CAE network consists of an encoder and a decoder, formulated as follows:
@@ III-D Behavior Cloning Network
We propose and design a BC network to learn the behaviors of coordinated manipulation skills of bimanual grasping from human demonstration data. Dexterous bimanual grasping skills can be considered into two categories: (1) adaptive interaction with objects, and (2) dual-arm motion coordination. To capture these skills, we have designed the input to our network to include encoded tactile feature maps, tactile image differences, and the robot’s proprioceptive state. The encoded feature maps and tactile image differences capture the human-object interaction skills. The robot’s proprioceptive state, on the other hand, offers insights into the coordination of movements between both arms. These inputs collectively serve to reflect the complexity and adaptability of dexterous grasping skills.
@@
Using such data of multiple modalities as input to train a network requires a well-crafted embedding structure [25]. A common way of fusing a 2D feature map and a 1D feature vector is to flatten the 2D feature map into a 1D vector and concatenate the flattened vector and the 1D feature vector [26]. However, we found that the flattening projection results in the loss of spatial correlation of tactile information. For fine dexterous pinch grasping of small objects, the spatial information provided by tactile images is essential in understanding the contact situation and the object being contacted. To preserve the spatial information of the encoded tactile feature maps, we specifically tile the proprioceptive state of robots and the tactile image difference to match the dimension of the tactile feature maps, so as to keep the spatial information of the encoded tactile feature maps.
@@
We then concatenate the tactile feature maps, the tiled proprioceptive state maps, and the tactile image difference on each feature channel. The convolutional layers in the BC network first filter the input feature maps (46$\times$16$\times$16) to a feature map (1$\times$8$\times$8), which is then flattened and fed into a fully connected network (FCN). The FCN network outputs a vector $\hat{\bm{a}}\in\mathbb{R}^{12}$ as the predicted Cartesian pose commands of the two arms, including 3D position and 3D orientation for each arm.
@@ IV-A Experimental Setup and Data Collection
We validate the performance of LfD with tactile sensing for robot dexterous manipulation on the challenging task: the retrieval of small and fragile objects from the desk using dual-arm pinch grasp and and ensuring a stable grasp throughout the process. During dexterous grasping, external vision can be easily occluded by the end-effector, potentially leading to inaccurate object estimation. Therefore, our experiments operate without using external visual sensors. By default, the starting position of the object lies between two robot hands, and the whole demonstration is operated in the task space.
@@
We collected 5 demonstrations for this task. The human demonstration dataset collected in the grasping task includes three main components: the Cartesian commands, the proprioceptive states, and the tactile feedback (i.e., tactile images provided by the TacTip sensors). The Cartesian commands and the proprioceptive states of the two arms are collected at a frequency of 1000 Hz. Two Tactips record the tactile image pairs at a frequency of 60 Hz. For each demonstration, about 1500 tactile images are recorded. Before using the collected dataset to train the networks, several pre-processing methods are used to process the raw data. The proprioceptive states of the two arms and the tactile images, collected at different sampling rates, are synchronized using a linear interpolation method to align their timestamps. A median filter is then applied to smooth the Cartesian commands $a_{t}$, i.e., the 6D poses of two end-effectors. For raw tactile images, the structural similarity index measure (SSIM) [27] is used to quantify the difference between the current frame and the original frame.
IV-B Implementation detail
Our proposed model is developed using PyTorch [28]. For the training of the CAE, we utilized a dataset that comprised 20 trials, including demonstrations of random behavior unrelated to our experiment. This dataset was collected using our robot setup, with each trial yielding approximately 1500 images captured during the demonstration phases. The trained CAE exhibits a satisfactory reconstruction quality, with a Mean Squared Error (MSE) loss of 0.015 and a Structural Similarity Index Measure (SSIM) of 0.934. The model training process for CAE, which involved 100 iterations, was completed in approximately two hours using an NVIDIA 1080 GPU. In the case of the BC network, the model was trained using data from 5 collected demonstrations, with the training process involving over 1,000 iterations and taking approximately 5 minutes to complete.
IV-C Design of Validation Tasks
IV-C1 Learning grasping vial
@@
The human demonstrator performs teleoperation of dual-arm robots to grasp a plastic vial (a test tube with $\Phi=15.65$mm) that is horizontally placed on the table. A Behaviour Cloning (BC) network is trained using the gathered demonstration data, and the trained policy is tested on dual-arm robots to validate its generalisation on unseen initial poses. During the evaluative phase, we positioned the test tube between the end-effectors to evaluate the performance of the learned policy given variations in the starting position, specifically alterations of up to ±$\pm 20$ degrees and displacements of up to ±$\pm 2$ centimetres in the objects’ locations.
IV-C2 Generalization to unseen objects
@@
IV-C3 Robustness against external disturbance
@@ IV-C4 Re-grasping capability
(b) Unsuccessful grasping of the baseline FCN policy trained with only the proprioperceptive information of end-effectors’ pose and positions.
Figure 4: Results of the comparison study. The policy trained with both comparison frameworks bypass the object directly, manoeuvring the end-effectors directly to the desired end-poses without making any physical contact, grasping attempts, or interactions with the tube.